Please make sure to double check your work for the correct answer & please show all steps! Thank you.For the most recent year, LMN, Inc., had sales of $467396, cost of goods sold of $234116, depreciation expense of $51178, and additions to retained earnings of $69054. The firm currently has 18455 shares of common stock outstanding, and the previous years dividends per share were $1.31. Assuming a 27 percent income tax rate, what was the times interest earned ratio? Answer to two decimal places.
Answer #1

Answer :

Revenue =467396

Cost of good sold,=234116

EBITDA =467396-234116=233280

EBIT =EBITDA - depreciation =233280-51178=182102

Let X be the interest payments

PBT =EBIT -x=182102-x

PAT = PBT *(1-TAX RATE) =(182102-X)*(1-27%)=132934.46-0.73X

Dividend= 1.31*total share =1.31*18455=24176.05

Retained earning=69054

PAT=24176.05+69054

132934.46-0.73X=93230.05

X=54389.60

Times interest earned ratio = EBITDA /interest payment =233280/54389.60=4.289